OpenStudy (anonymous):

What is the angular speed (in rev/min) of the Earth as it orbits about the Sun?

OpenStudy (unklerhaukus):

1 revolution per year

OpenStudy (turingtest):

\[\frac{2\pi\text{radians}}{\text{revolution}}\times\frac{\text{revolution}}{\text{year}}\times\frac{\text{year}}{365\text{days}}\times\frac{\text{day}}{24\text{hours}}\times\frac{\text{hour}}{60\text{minutes}}\]

OpenStudy (turingtest):

or\[\frac{\text{revolution}}{\text{year}}\times\frac{\text{year}}{365\text{days}}\times\frac{\text{day}}{24\text{hours}}\times\frac{\text{hour}}{60\text{minutes}}\]since you just want it in terms of revolutions...

OpenStudy (unklerhaukus):

angular speed is \[1.90 \times 10 ^{-6} [{rev \over min}]\]